Carnegie Mellon Robotics Institute | NASA ASTEP | EventScope

Life in the Atacama Header

Home

News Down Arrow
 • Field Reports
 • Media Articles

Partners

Information Down Arrow

Astrobiology Down Arrow

Robotics Down Arrow

Images Down Arrow

Movies

Science Data
(best viewed in Firefox)

Workshops

Add Yourself to the Atacama Mailing List


View the Atacama in 3D with Eventscope Button

September 9-10, 2004
Salar Grande, Atacama Desert, Chile

Agenda
Finalize fluorescence imaging protocol
Test onboard spectrometer with foreoptic
Work around failed power relays
Drive autonomously

Status and Progress
Fluorescence protocol developed. We continue to have excellent results from Zos onboard fluorescence imager (FI). After initial confirmation that the instrument is able to excite and detect fluorescence in daylight (under the shade of the rover), we have further refined the specific filter combinations and exposure times in order to reliably detect chlorophyll fluorescence with 450nm excitation working better than 350nm. Tests with sprayed water we found to significantly improve the fluorescence signal of chlorophyll containing organisms, perhaps simply due to removing dust. A DNA dye, which only fluoresces in the presence of DNA, was found to provide minimal enhancement perhaps due to the difficulty in penetrating the exterior coating of microorganisms.

Onboard spectrometer functioning. We have verified the onboard visible-near infrared (VNIR) (350nm-2500nm) is functioning properly through the 1 foreoptic mounted on Zos pan-tilt mechanism. The pan axis makes a complete rotation, 360 and the tilt axis covers the sky (+20) to almost straight down (-80). This configuration allows Zo to collect collect images but also spectra of targets anywhere in its vicinity. The interface electronics between the spectometer and Zos computers become more stable as we refine the boot sequence, and as the protocol we use for collecting spectra gets more concrete we are better able to keep the device in proper dynamic calibration, specifically to maintain proper dark current measurements and white references (for changing light levels). A goal of this years field experiments is to get sufficient experience with rover-based VNIR to reliably collect spectra and have a working protocol to use in the future.

Replaced power relays. Further diagnosis will be required to understand exactly why Zos main power relays are slowly failing, but a likely explanation is that the are unable to tolerate low voltage conditions that exist during low-power startup. The relays have an integral circuit that sets the voltage at which the relays pull in and that drops the current required to hold them, it seems that it is this circuit, not the relay itself that is failing. Our work around is a manual switch to turn off and on rover power. The implication is that Zo will not be able to turn itself off and on.

Drove autonomously. We have conducted some initial tests in autonomous navigation, in the vicinity of base camp, to tune the camera models and traversabililty estimates. As expected, late day low-angle sun conditions have caused some difficulties but performance is generally good. In the coming days we will move to longer traverses out of camp.

Upcoming
Exercise autonomous navigation
Continue to refine position estimation
Test autonomous mission planning and execution
Prepare for first remote science experiments

Weather
Mornings: Light fog, still, moderate 11-13C, 70% humidity
Afternoons: Clear, some light breezes, warm 23-25C, 40% humidity
Evenings: Low fog, calm, warm 14C, 50-60% humidity

Quote of the Day
"We continue to enjoy spam in the Atacama"
"guru chunky whomsoever francine additional brouhaha"

For more information on Life in the Atacama including images, movies, and field reports see: http://www.frc.ri.cmu.edu/atacama
Logos

© 2003-2008 Carnegie Mellon University. All rights reserved. Webmaster.

Carnegie Mellon Logo Robotics Institute Logo Eventscope Logo NASA Logo